# Wave-induced drift in third-order deep–water theory

### v1

@@ 5.3 Random waves
Because we are interested in comparing different computations of drift, we also derive a generalised version of the Stokes drift (42) with second order difference harmonics suitable for arbitrarily many modes. The procedure is familiar: Taylor expand the velocity field from (22) about an initial position , and keep the lowest order terms. Solve for the particle trajectories by integrating that expanded velocity field in time. Then substitute the solution thus obtained into the second order terms in the Taylor expansion, and integrate the result to find the corrected particle trajectories. We obtain
Here we recognise term (I) as simply the sum of the individual Stokes drifts, and term (II) as the generalisation of (42).
@@
Figure 14 shows the results of computing the wave-induced particle displacement for random waves. The amplitude spectrum (45) is retained, but rather than assigning the phases of the waves so as to achieve linear focusing at a specific location, these phases are randomly generated for 200 realisations. Red curves denote first order theory, green curves second order, and blue curves third order, with the prominent, dark curves representing the average over all realisations. In each case the integration is for approximately 80 peak periods, which is long enough to clearly distinguish the particle drift.
As we observed for focused waves in Figure 13 and bichromatic waves in Figure 10, near the surface the linear theory gives the largest forward drift. While particular realisations (shown in lighter curves) may exhibit a larger drift in a nonlinear (2nd or 3rd order) theory, on average the linear theory is an overestimate of the forward drift. Likewise, as was the case for focused waves and bichromatic waves, the Stokes drift formulations represent an underestimate of the drift. Unlike specific cases identified for bichromatic waves in Figures 7 and 10 there is no discernible difference in the Stokes drift when difference terms (II) are included near the surface.
@@
The more interesting setting for these random waves is observed at greater depths, which is shown in panel (b). With a peak wavenumber 1/m, most of the motion due to the free-wave constituents is filtered out at a depth of m. The linear theory gives a very narrow range of displacements, with the realisations in red barely showing deviation from the average, and agreeing well with the linear Stokes drift (I). Second and third order displacements are quite large in certain realisations, depending on the phase relationships involved. However, as expected, there is no systematic return flow at any depth for random waves, because such (spatially and temporally) localised flows are averaged out to yield the net forward drift which we observe in Figure 14(b). Moreover, all theories approach zero from above with depth, exhibiting the same ordering we have hitherto observed: linear theory predicting the smallest drift followed by second and third order. The correction (II) in (46) is found to give a useful improvement over using only the classical Stokes drift term (I).
6 Conclusions
@@
The wave induced drift of particles has attracted increasing attention in recent years, particularly in connection with the transport of maritime pollution. Nevertheless, fundamental explorations of this topic are rare. Notable exceptions are the recent articles by van den Bremer & Taylor (2016) and the follow-up study by van den Bremer et al. (2019), which focused on the transport associated with wave groups to second order. Our aim has been to explore particle drift systematically using the first, second, and third order theory, and working our way from monochromatic waves to random seas. This work rests on numerical integration of the particle trajectory mappings and comparison with simplified formulas like that for Stokes drift.
We consistently observe that linear theory gives the largest forward drift close to the free surface. Adding second order effects decreases this drift somewhat when more than one mode is present. The reason is simple: where the linear waves have the largest displacement we find a region of negative velocity in the second order difference harmonics. Third order effects decrease the forward drift further through a combination of changes to wave speed – which shrinks the particle paths – and the addition of further harmonics.
@@
Descending in the water column we find a different overall behaviour of the wave-induced drift. It is well known that the water column acts like a filter, first suppressing the highest frequencies and filtering out ever lower frequencies with increasing depth. Particle motion is typically considered negligible half a wavelength below the surface. It is this filtering effect on linear, free-wave constituents that leads to the dramatic decrease of drift in the linear theory. However, nonlinearity introduces much longer bound-waves which become dominant at depth. Beneath the centre of a wave group these can induce sizable localised return flows. However, their overall contribution is to increase the drift in the direction of wave propagation versus linear theory. Third-order contributions lead to additional bound waves, but also to dispersion corrections; the overall effect of third-order vs. second-order nonlinearity thus depends on parameters of the sea-state such as its bandwidth and amplitude spectrum.
These results for each order require integration of the particle trajectory mapping, which is inconvenient. The more common approach in practice is to use the Stokes drift formula, which approximates the linear theory via a Taylor series expansion. Comparisons show that this approximation is quite poor, particularly close to the surface, where linear theory yields a significantly larger drift. In fact, the Stokes drift formulation inadvertently lies quite close to the smaller drift predictions of the nonlinear (second or third order) theories near the surface, which may explain some of its success.
At depth the Stokes drift hews close to the linear drift, but we have seen that this may differ significantly from the drift in nonlinear wave theories. To this end, we have retained a further term in the Stokes drift formulation which bridges some of this gap, and accounts for second order difference harmonics. For settings with multiple nearby harmonics this adjusted Stokes drift formula captures the larger forward drift associated with nonlinear waves, and represents an improvement on the commonly used formulation.
@@
Our objective has been to provide a systematic and transparent exposition of drift associated with waves in nonlinear potential flow. To this end numerous simplifications have been made. Noteworthy among these is the neglect of viscous effects at the free surface, which has been known since work of Longuet-Higgins (1953) to play an important role. We have likewise neglected directional effects, which can be expected to significantly complicate any study. Finally, and significantly, we have relied on the simplification of infinite water depth. This has allowed us to present numerous simple and compact formulations which are difficult to find in the existing literature, and contributed – we hope – to a clear presentation. Nevertheless, it is important to note that long bound harmonics may “feel the bottom” even when the free waves do not, and so the inclusion of finite depth is of central importance. This will be treated in a subsequent work.
With all of the caveats above, it is important to mention that the drift incorporated into operational models such as WAVEWATCH III is the simplest infinite-depth Stokes drift formulation (Li et al., 2016). This points to a continued need to better understand this formulation – with all the simplifications it involves – and how it may be further refined. For example, it has been pointed out recently (Lenain & Pizzo, 2020) that correct treatment of the high-frequencies is critical for estimating the surface drift. The high-frequency components of the wave spectrum undergo the largest correction due to third-order effects (Stuhlmeier & Stiassnie, 2019), and further investigation of how to correct the Stokes drift to account for this could be highly relevant for practical applications. It may also be of interest to go beyond the confines of the third-order theory presented here, and to explore particle trajectories for solutions to the full Euler equations e.g. via conformal mapping techniques.
@@ Appendix A On the non-uniqueness of Stokes expansion
The former (47)–(48) is given by (Wehausen & Laitone, 1960, (27.25)), and appears naturally when the third order potential is set to zero. However, at each order solutions to the Laplace equation permit the addition of first-harmonic terms of the form . The latter (49)–(50) is the natural form obtained from the Zakharov formulation (see also (Janssen, 2009, (A.16)), (Gao et al., 2021, (6.7c))). Note that this flexibility in the perturbation expansion explains the seeming discrepancy identified by Zhang & Chen (1999) when matching these solutions.

### v2

@@ 5.2 Random multichromatic waves
Where focused wave groups exhibit localised features, such as a clear division between flows in the direction of propagation at the surface and return flows at depth, – which have been well described by using wave-envelope formulations by van den Bremer and coworkers (Van Den Bremer & Taylor, 2016; van den Bremer & Breivik, 2018; van den Bremer et al., 2019; Calvert et al., 2019) – the lack of clear “groupiness” in random seas means that we should expect these localised effects to be averaged out. The advantage is that we can test whether (on average) the Stokes drift formulations developed in (41) are suitable, since they clearly cannot capture spatio-temporally localised flows around a group itself (all terms are positive).
@@
Figure 13 shows the results of computing the wave-induced particle displacement for 200 realisations of a random sea surface. The amplitude spectrum (40) is retained, and the same parameters 1/m and as in Figure 11 are used, but rather than assigning the phases so as to achieve linear focusing at a specific location, these are randomly generated for each realisation. Blue curves denote first order theory, red curves second order, and yellow curves third order, with the prominent, darker curves representing the average over all realisations for a given order. In each case the integration is for approximately 80 peak periods, which is long enough to clearly distinguish the particle drift.
@@
With a peak wavenumber 1/m, most of the motion due to the free-wave constituents is filtered out at a depth of m. The linear theory gives a very narrow range of displacements, with the realisations in red barely showing deviation from the average, and agreeing well with the linear Stokes drift (SD I). Second and third order displacements are quite large in some realisations, depending on the phase relationships involved. However, as expected, there is no systematic return flow at any depth for random waves, because such (spatially and temporally) localised flows are averaged out to yield the net forward drift which we observe in Figure 13. Moreover, all theories approach zero from above with depth, exhibiting the same ordering we have hitherto observed: linear theory predicting the smallest drift followed by second and third order. The correction (II) in (41) is found to give a useful improvement over using only the classical Stokes drift term (I).
5.3 Parametric wave spectra
The comparisons of the previous sections, based on explicit computation at different orders, support the inclusion of quartic difference harmonic terms in calculations of the Stokes drift throughout the water column. For many applications in ocean science and engineering it is advantageous to present such a result in terms of energy rather than amplitude spectra, and to explore its consequences for some common parametric spectral shapes.
For a continuous wavenumber spectrum , we make the identification with the modal amplitudes where the uniform grid spacing is assumed to tend to zero. With this assumption, is is straightforward to rewrite the discrete multi-chromatic Stokes drift with second order contributions as
where the first term is identical to that found by Kenyon (1969) and the second term comes from the inclusion of second-order terms in deep water.
Numerous critical discussions have focused on the key role of the high-frequency cut-off on the Stokes drift near the surface, including Breivik et al. (2014) and more recent work by Clarke & Van Gorder (2018) and Lenain & Pizzo (2020). As Clarke and van Gorder (2018) point out, if the Stokes drift is rewritten in terms of frequency and the high-frequency tail of the spectrum is such that , then we have a divergent expression for . Their proposed resolution, introducing a wave breaking frequency as an upper limit for spectral Stokes drift calculations was found to be reasonable for the subsurface Stokes drift by Lenain & Pizzo (2020), but we will employ the maximum wavenumber suggested by Lenain & Melville (2017) and used by Lenain & Pizzo (2020).
As a first comparison, we consider the Phillips spectrum for the equilibrium range of wind-generated waves, given in terms of radian frequency as
where we take as in Breivik et al. (2014). We relate the peak frequency to a friction velocity (see Holthuijsen (2007)) in order to establish as above, and show the results in Figure 14.
Panel (a) shows the Stokes drift near the surface using either the first term in (42) only (SD I) or both terms in the same equation (SD I+II). The difference between the two is sizable, but strongly dependent on the choice of high-frequency cut-off This tracks with the very pertinent discussion in Lenain & Pizzo (2020) (therein they report an underestimation of the Stokes drift of up to 50% from failure to account for high frequencies, but they do not consider the effect of difference harmonics). This difference persists as we descend into the fluid, and a noticeable difference on the order of 1 % is visible at depths below one peak wavelength in panel (b). Defining the Stokes transport as
we find nearly a 9% change in total Stokes transport. If we calculate the Stokes transport only below of the -folding depth of the peak frequency, we still observe a change of approximately 2% owing to the inclusion of difference harmonic terms (II).
The Phillips spectrum describes only the high-frequency tail, so it is of interest to consider a more realistic parametric spectrum: we next consider a Pierson-Moskowitz (PM) spectrum, corresponding to a fully developed under a constant wind over unlimited fetch. The case shown in Figure 15 corresponds to a wind speeds of 10 m/s (with corresponding ), and peak frequency of 0.17 Hz. Blue curves show the formula of Kenyon (1969) (corresponding to the first term in (42), labelled SD I) and red curves show both terms of (42) (labelled SD I+II). For reference, we note that the high-frequency cut-off obtained from is 6.3 We note also that it is possible to cut off the low frequencies, and evaluate the outer integrals in (42) starting from instead of 0 without materially affecting the results, as suggested by Breivik et al. (2016); Lenain & Pizzo (2020).
The differences in the formulations are clearly visible, both near the surface and at depth. At the surface, we find increases of surface drift in excess of 40% when difference harmonic terms are accounted for, reducing to just less than 2% at half the -folding depth of the peak wavenumber Interestingly, starker differences reappear at greater depths, as shown in panel (b) which highlights the different asymptotics of the two formulations at depths of more than one peak wavelength. We find that keeping both terms in (42) leads to an increase in Stokes transport of nearly 3.3 % for m/s. The majority of this difference comes from the near-surface transport, with transport below the peak -folding depth accounting for a difference of only 0.1%. The same analysis can be carried out for JONSWAP spectral shapes, with qualitatively very similar results to those found in Figure 15.
We can also compare the Stokes drift formulations of equation (42) for other spectral shapes, such as swell-wave spectra described by Ochi & Hubble (1976). In Figure 16 we show the most probable Ochi-Hubble shape corresponding to a significant wave-height m, and the two Stokes drift calculations. The surface drift shows a change in excess of 20 % when subharmonic terms (II) are included, and the Stokes transport is slightly more than 1% larger over the entire water column.
6 Discussion and conclusions
@@
The wave induced drift of particles has attracted increasing attention in recent years, particularly in connection with the transport of maritime pollution. Nevertheless, explorations of the underlying hydrodynamics are not widespread. Notable exceptions are the recent articles by Van Den Bremer & Taylor (2016) and the follow-up study by van den Bremer et al. (2019), which focused on the transport associated with wave groups to second order. Our aim has been to explore particle drift systematically using the first, second, and third order theory, working our way from monochromatic waves to random seas.
Monochromatic waves – or progressive Stokes waves – are extremely well studied, and therefore serve as a useful test case for any development of nonlinear wave theory. They allow us to compare with exact solutions at the surface (Longuet-Higgins, 1987) as well as high-order approximate solutions in Eulerian (Zhao & Liu, 2022) and Lagrangian (Clamond, 2007) variables. The ease of defining a monochromatic wave – in the absence of an Eulerian current, which is the case for all the waves we consider, only an amplitude and a period are needed (Fenton, 1990) – means that approaches at each order can be directly compared to one another. Such comparisons show that the classical Stokes drift formula is remarkably accurate, at least when compared to other inviscid, irrotational theoretical results. Some differences do emerge for waves of large steepness, where the Stokes drift formula derived from linear wave theory tends to underestimate drift at the surface and overestimate it at depth.
With a view towards understanding more complex patterns of waves (and wave groups) on the open sea, we have subsequently studied the unsteady flow associated with bichromatic waves, generated at lowest order by two Fourier harmonics. Unfortunately, very few experiments have been carried out on bichromatic waves, and available results such as those by Westhuis et al. (2001) focus on the free surface elevation and the development of instabilities rather than induced flows. Other comparisons of horizontal velocities in bichromatic waves (Stansberg et al., 2008) rely on numerical simulations only. The problem is compounded by the fact that bichromatic waves are not themselves well-defined, since the inclusion of higher-order effects leads to a phase shift between the modes. This makes comparisons between different theories fraught with difficulty, and despite the fact that the bichromatic wave train is an exact solution to the 3rd order water wave problem Mei et al. (2018) no exact numerical solutions and attendant kinematics appear to be available in the literature.
@@
That said, our investigations are able to shed some light on the flows induced by bichromatic waves. It is well known that the water column acts like a filter, first suppressing the highest frequencies and filtering out ever lower frequencies with increasing depth. Therefore the particle motion of monochromatic waves is typically considered negligible half a wavelength below the surface. It is this filtering effect on linear, free-wave constituents that leads to the dramatic decrease of drift in the linear theory for bichromatic waves. However, nonlinearity introduces much longer bound-waves which become dominant at depth, as was recognised in the landmark work of Longuet-Higgins & Stewart (1964). Beneath the centre of a wave group we find that these can induce sizeable localised return flows. However, their overall contribution is to increase the drift in the direction of wave propagation versus linear theory. Third-order contributions lead to additional bound waves, but also to dispersion corrections, which tend to retard the drift slightly.
In the bichromatic case the potential at second order includes new – difference – harmonics, in contrast to the monochromatic potential, which is identical at first and second order. This provides an impetus to update the Stokes drift formula, simply by using the second order potential in the Taylor series expansion in place of the first order. The new term – which captures the Stokes drift of second-order difference harmonics, can make a substantial contribution to the depth dependent drift depending on the wavenumbers involved. This contribution is also captured when simply integrating the particle trajectory ODEs in second- or third-order theory, as well as when using a fourth order Lagrangian drift due to Blaser et al. (2025). This corroborates the idea that modifying the Stokes drift formula by the inclusion of difference harmonics is a good idea, with the important caveat that all modes (including subharmonics) are assumed in deep water (see below).
Waves with more than two fundamental harmonics at lowest order become increasingly complex - from an algebraic perspective when deriving an analytical theory, and from a practical perspective when attempting to meaningfully describe the various parameter regimes. As with bichromatic waves, for focused wave groups we see strong localised displacements at depth in the opposite of the wave propagation direction, but find that the averaged drift remains positive, with slightly larger values for nonlinear theories that include the effects of subharmonics. We explore the consequences of extending the ubiquitous formulation of Kenyon (1969) by including difference harmonic terms for energy spectra, and obtain significant changes to surface Stokes drift and Stokes transport for several parametric spectra. These changes are strongly dependent on resolving high frequencies (see Lenain & Pizzo (2020)), and it would be interesting to see if such effects can be seen in field experiments or suitable ensemble-averaged flume experiments.
@@
Our objective has been to provide a systematic and transparent exposition of drift associated with waves in nonlinear potential flow. To this end numerous simplifications have been made. Noteworthy among these is the neglect of viscous effects at the free surface, which has been known since work of Longuet-Higgins (1953) to play an important role. We have likewise neglected directional effects. Although our methodology can be adapted without change to include wave directional spreading, the additional freedom that this provides makes for a significant additional burden. Nevertheless, directional spreading is key when comparing with real sea states, and it would be of interest to follow in the footsteps of the recent work by Higgins et al. (2020) and explore these effects. Finally, and significantly, we have relied on the simplification of infinite water depth. This has allowed us to present numerous simple and compact formulations which are difficult to find in the existing literature, and contributed – we hope – to a clear presentation. Nevertheless, it is important to note that long bound harmonics may “feel the bottom” even when the free waves do not, and so the inclusion of finite depth is of central importance. Indeed, finite depth effects can also notably shape the particle drift for monochromatic waves, a subject which has been the subject of exploration from the work of Longuet-Higgins (1953) to the present day (Grue & Kolaas, 2017). We intend to treat finite depth waves in a subsequent work.
There are also numerous other cases of interest that should be mentioned. It might be argued that, after Stokes waves the next most-complex – and already unsteady – solution is that of standing waves. While some higher-order solutions are available in both Eulerian (Bryant & Stiassnie, 1994; Schwartz & Whitney, 1981) and Lagrangian (Chen & Hsu, 2009) variables, the theory is not nearly as complete as that for periodic, progressive waves. A systematic exploration of the kinematics of standing waves at various orders, in line with that suggested here for progressive waves, would be an interesting area for future work. Because of our assumptions of infinite water depth for all modes and a vanishing Eulerian mean flow it is quite difficult to compare with experimental results obtained in closed wave flumes, wherein Eulerian mean flows are typically depth-dependent, and therefore not irrotational. However, our methodology could be adapted to include uniform currents, and so enable a comparison of higher-order Eulerian and Lagrangian kinematics (the latter derived by Chen et al. (2012)) in this setting.
The fact that a weakly nonlinear theories in Eulerian variables generally make use of Taylor expansion to transfer the boundary conditions to a half-space (and therefore confine the fluid domain to ) has unfortunate consequences. It has given rise to a significant literature aimed to deal with the problem of evaluating kinematics in the surface zone between the crest and trough level, and particularly at wave crests, beginning with the stretching methods of Wheeler (1970) and continuing to the present day. Recent comparisons of these methods can be found, for example, in Stansberg et al (2008) or Johannessen (2010). We have not resolved the problem of identifying kinematics between crest and trough in Eulerian variables, but we have attempted a theoretically consistent, order-by-order exploration of a variety of waves.
With all of the caveats above, it is important to mention that the drift incorporated into operational models such as WAVEWATCH III is the simplest infinite-depth Stokes drift formulation (Li et al., 2016). This points to a continued need to better understand this formulation – with all the simplifications it involves – and how it may be further refined. We hope that the present work presents a modest step in this direction.
@@ Appendix A On the non-uniqueness of Stokes expansion
The former (44a)–(44b) is given by (Wehausen & Laitone, 1960, (27.25)), and appears naturally when the third order potential is set to zero. However, at each order solutions to the Laplace equation permit the addition of first-harmonic terms of the form . The latter (45a)–(45b) is the natural form obtained from the Zakharov formulation (see also (Janssen, 2009, (A.16)), (Gao et al., 2021, (6.7c))). Note that this flexibility in the perturbation expansion explains the seeming discrepancy identified by Zhang & Chen (1999) when matching these solutions.
When dealing with monochromatic waves, it is important to note that the solutions (44a)–(44b) and (45a)–(45b) need to be adjusted when a wave of fixed height is considered. According to these formulations, the height of the wave, defined as is related to the linear amplitude by either
where subscript 1 refers to (44a)–(44b) and subscript 2 to (45a)–(45b). For a prescribed height these expressions can be iteratively inverted to yield
Inserting this into the third order solution recovers the coefficients in terms of found, for example, in Fenton Fenton (1990).
@@ Appendix B Coefficients of the higher-order potential
Appendix C Effects of amplitude evolution on particle kinematics
In Section 2.2 we contend that it is appropriate to neglect the evolution of Fourier amplitudes while calculating particle trajectories and associated drift. This is based on the scale separation between the slow amplitude evolution (taking place on a time-scale of ) and the much faster particle motions, whose time-scale is in principle , i.e. proportional to the wave period. As we have seen, the motion of an individual particle depends on the wave phase, and so a single estimate of Stokes drift is – in some sense – an average value only. This is borne out by our methodology for calculating the horizontal drift of particles using 1st, 2nd, and 3rd order theory for monochromatic waves (Figure 4), bichromatic waves (Figure 8) and multichromatic waves (Figure 13).
We explore the role of amplitude evolution for an extreme example in Figure 17, which considers the evolution of a modulationally unstable wave with 1/m and (linear) steepness together with two initially small disturbances and for For clarity, our comparison excludes 2nd and 3rd order bound modes (see Sec. 2.3), and focuses only on the effects of amplitude evolution.
At the initial time the free surface is essentially monochromatic (Fig. 17(b)) and nearly all of the energy is the carrier wave . Subsequently, instability gives rise to energy exchange between the carrier and the side-bands and which can be captured by solving the Zakharov equation (8), and is shown in Figure 17(a). Note that the characteristic time for energy exchange, s, where is the carrier period.
Panel (c) of Figure 17 shows a particle trajectory at an initial depth of m. The particle is tracked for 10 starting at time s – indicated by black curves on the left of panel (a). The amplitude evolution here is such that the energy is being transferred from the carrier to the side-bands, until equipartition is reached in the Fourier amplitudes. The energy transfer is captured in Panel (d), which shows the initial Fourier amplitudes in red (carrier larger than side bands) and the final Fourier amplitudes after 10 in blue. In Panel (c) we compute the particle trajectory using both the initial amplitudes only (red curves) and with time-dependent amplitude evolution of Panel (a) (blue curves). Panel (e) shows the corresponding evolution away from equipartition in amplitudes, i.e. starting at as shown in Panel (f), and following the amplitude evolution for 10 along the black curves in Panel (a).
As alluded to earlier, this is an extreme case characterised by a very narrow spectrum and a dramatic spectral evolution from monochromatic waves to a strongly modulated wave train (see Panel (b)). Nevertheless, the amplitude evolution is essentially negligible over a single carrier period , and not dramatic even over time-scales of Moreover, any change in particle paths due to amplitude evolution depends both on the phase-position of the particle as well as where we are in the cycle of energy exchange.
In Figure 18 we consider a rather milder type of nonlinear interaction (though still significant enough to be interesting), where a Gaussian spectrum (with 1/m, 1/m, and m as per (40)) consisting of 10 Fourier modes with maximum initial slope is allowed to evolve according to the cubic Zakharov equation (8). The energy exchange between modes leads to changes in the packet-structure of the free surface (shown in Panel (a)), owing to the spectral broadening which is illustrated in Panel (e), which is solved for approximately .
Panels (b) and (c) of Figure 18 show particle paths evolving for 10 peak periods starting at and s, respectively. As in Figure 17, blue curves include the amplitude evolution captured by the Zakharov equation while red curves employ constant, initial amplitudes. Unsurprisingly, given the slow rate of spectral change, close to s there is scarce difference in the particle paths over 10 in panel (b). On the other hand, further into the time evolution the wave fields are clearly different, and particle paths cannot be expected to agree. Despite this, when averaging the drift of particle trajectories over 20 peak wavelengths we find the Stokes drift for evolving amplitudes agrees very well with the result for constant amplitudes, shown in Panel (d). In fact, due to the exclusion of bound modes we find that the Stokes drift is very close to the linear formula (I) from (41) – shown in yellow circles in panel (d).
Recent work by Blaser et al. (2025) studied the Stokes drift of wave packets undergoing both dispersive focusing and amplitude evolution using a fully nonlinear potential flow solver. Our focus in this appendix is somewhat different from theirs, as we do not concentrate on an isolated wave packet, nor examine the transport for the entire packet; nevertheless, the extent to which a simple cubically nonlinear theory such as that developed here could explain some of the findings from their numerical simulations remains an interesting question for future work.